# Global Streams, Local Currents: A Data Analysis on Global VOD Content Consumption

### v1

@@ Abstract
This study explores global VOD (video on demand) content consumption patterns through a network-based approach. We used Netflix’s ’TV-shows’ ranking data, spanning 822 days across 71 countries, to construct a network where countries are represented as nodes and consumption similarities are reflected as link weights. By applying the Louvain algorithm, we identified three distinct consumption groups: ‘North America and Pan-Europe’, ‘Asia and Middle East’, and ‘Central and South America group’. These groups align closely with geographic, historical, and linguistic divisions, despite no predefined grouping criteria. Notably, Turkiye, often considered a cultural and regional crossroads, exhibited some classification ambiguity but was ultimately grouped with Asia and Middle East. Our findings also reveal the dominant influence of the United States across all groups, while South Korean content, particularly after the success of “Squid Game” in 2021, has gained and sustained popularity in Asia and Middle East and Latin America. This study, based on data, demonstrates that deep-seated cultural histories continue to shape global consumption patterns, even amidst rapid changes in media platforms and content production dynamics.
@@ 1 Introduction
Despite of the de-globalization trends in realm of politics and economy, culture is in an unprecedented state of globalization. VOD (Video on Demand) services such as Netflix, are the primary drivers of transnational popular culture. As seen in traditional global media consumption patterns based on cultural proximity[31] however, the consumption patterns of VOD services may heavily depend on geographic proximity. Indeed, Netflix has managed their market by dividing it into four geographic cultural clusters (Europe-Middle East-Africa, Asia-Pacific, The United States-Canada, Latin America).
@@
Despite interest in VOD acceptance[32], empirical research became possible only after Netflix’s TOP 10 consumption data was made public in 2020. Lotz[22] has analyzed the frequency of overlap in the preferred Netflix Top 10 titles by major countries and argued that global consumption of Netflix consists of large consistent markets such as English-speaking and Spanish-speaking regions, and small variable markets such as geographically close Korea, Japan, Taiwan, and the unique market of India. Jang et al.[15] have identified eight groups that consume Netflix similarly based on Netflix’s Top 10 consumption data, discovering that they are differentiated by geographical distance, linguistic similarity, and cultural differences.
@@
The concept of geo-cultural or geo-linguistic markets shows that global VOD services are consumed differently regionally, geographically, and linguistically. According to their diversity report[26] in 2021, Netflix places a strong emphasis on producing content that reflects diverse regions and cultures, in addition to being inclusive in terms of race and gender. The strategy is described as a deliberate response to engage with and cater to unique audiences in different regions[6]. While the geo-cultural and geo-linguistic market concepts do demonstrate that global VOD services are consumed differently regionally, geographically, and linguistically, there remain questions. These include inquiries into how global VOD consumption is defined and separated, which nations are pivotal in this segmentation, and beyond it. It is also essential to persistently monitor the evolving patterns of consumption among similar viewer groups over time with longitudinal data. What distinct viewer groups emerge from the complex interactions of people connected to VOD platforms worldwide?
@@
This study aims to address the question by exploring how global VOD is segmented into different national and regional markets. These divisions may not just by country borders, but also by larger cultural and linguistic areas which can transcend national boundaries. Essentially, we examine how the consumption of VOD services is influenced by and caters to these diverse geo-cultural segments. To give a preliminary report, we found that global VOD content is being consumed in three distinct geo-linguistic consumer groups worldwide. Interestingly, these groups follow the trajectory of human history from the West to the East, from settled to migratory communities (alternatively, from colonizers to the colonized). Some countries are located on the boundary of these consumption groups, with Turkiye being a prime example. Due to its unique historical and geographical location straddling the East and West, Turkiye exemplifies a nation that could fall into more than one of these consumption categories. These patterns are not only geographical but also historical, reflecting human movement and cultural development from ancient to modern times.
This suggests that the ways in which global audiences choose and consume VOD content may be influenced by deep-seated historical and cultural trends. The consumption trajectory of global VOD services, which audiences themselves choose according to their preferences, represents ‘the long future’ predetermined by lengthy human history. ‘The long future’ calls to mind “The Long Revolution”, a monumental work by Raymond Williams[33], author of “Television: Technology and Cultural Form” in 1974[27]. Williams elucidates that unlike the Industrial or the Civil Revolutions, the cultural revolution was a lengthy journey that culminated with the emergence of mass media in the 20th century, long after the spread of literacy and the emergence of a reading public. The term “long future” is used to imply that the historical evolution of human society, including such milestones as the cultural revolution Williams discusses, continues to exert a significant influence on current global trends in VOD media consumption. In other words, the extensive path of our collective past continues to shape how we consume media today, and it suggests that these historical influences will persist as a powerful force shaping media consumption well into the future. Thus, we could call it global streams, local currents. The global VOD consumption patterns shown in this study are not hasty conclusions drawn from a few pieces of data, but rather optimized results discovered through more than a thousand computer simulations of actual VOD consumption data.
@@ 2.1 Data
We analyzed content lists of NS rankings, spanning from the 1st to the 10th place, collected from Flixpatrol over a period of 822 days ($N_{day}$) [^1] July 1, 2020 to September 30, 2022 and across 71 countries ($N_{coun}$) [^2] The 71 countries mentioned are the sum of the members of the three groups listed in footnote 4.. If each content appeared in the list only once across the period, the total number of content entries would be $822\times 71\times 10=583,620$. However, since most content appears in the list across multiple days and in different countries, the actual number of unique content entries was reduced to $N_{cot}=1769$.
@@
$G_{A}=\{g_{A}^{1,1},g_{A}^{1,2},\ldots,g_{A}^{2,1},\ldots,g_{A}^{822,1769}\}$ (1)
@@
As equation (2) indicates, the 1st place ranking is converted to 10 points, the 10th place to 1 point, and rankings beyond the 10th place are assigned 0 points. By converting ranking data into score data, we can now represent each country as a point in a high-dimensional space of $822\times 1769=1,454,118$ dimensions. The coordinates of a country in this space are given by its score data. For example, the position of country A in this hyperspace is represented by the set of scores, $S_{A}=\{s_{A}^{1,1},s_{A}^{1,2},\ldots,s_{A}^{2,1},\ldots,s_{A}^{822,1769}\}$. Within this coordinate system, we define the distance between two countries as the Euclidean distance in hyperspace, as detailed below.
@@ 2.2 NSC distance between countries
We aim to define the ‘NSC distance’ between two countries as the difference in their content consumption patterns throughout the entire analysis period. However, due to heuristic reasons, we first define the daily distance ($dd_{AB,d}$) between two countries, A and B on day d based on the differences in the ranked contents between two countries as follows,
$dd_{AB,d}=\sqrt{\sum_{c=1}^{N_{cot}}{\left(s_{A}^{d,c}-s_{B}^{d,c}\right)^{2}}}$ (3)
@@
In the summation, the terms $\left(s_{A}^{d,c}-s_{B}^{d,c}\right)^{2}$ calculate the squared difference in scores between countries A and B for each piece of content c on day d. This approach ensures that all differences are positive and gives more weight to larger discrepancies, regardless of which country had the higher score. Previous research[15] has focused on whether contents were in the top 10 list or not, without considering their specific rankings. We argue that taking ranking scores into account is crucial because the disparity between the 1st and 10th ranks is often more significant than between listed and unlisted content. This is particularly true given the power-law distribution observed in the popularity of cultural contents[3, 13, 25] which suggests that higher-ranked content has disproportionately greater popularity and influence.
@@
In equation (3), the summation index goes from 1 to $N_{cot}=1769$, but in practice, we only add up at most 20 terms that are not zero. For example, if content c is not in the top 10 for country A, then the corresponding term $s_{A}^{d,c}$ is zero, and the same rule applies to country B. This implies that the only terms you need to consider are those where the content is c is in the top 10 for country A or B. All other terms will be zero and can be ignored in the sum.
The NCS distance ($D_{A,B}$) between two countries, A and B, is defined as the square root of the sum of the squares of daily distances $dd{A}{B}{d}$ between them over the days in the entire analysis period in accordance with the Pythagorean theorem:
$D_{A,B}=\sqrt{\sum_{d=1}^{N_{day}}dd_{AB,d}^{2}}=\sqrt{\sum_{d=1}^{N_{day}}\sum_{c=1}^{N_{cot}}\left(s_{A}^{d,c}-s_{B}^{d,c}\right)^{2}}$ (4)
Some may argue that this distance may not accurately capture the differences in actual NSC patterns between two countries due to the limitations set by equation (4). Specifically, one concern arises because the calculation considers only content ranked in the top 10, excluding any data on ranks below the 10th, which could potentially distort the measurement of consumption differences. However, we believe this limitation is negligible due to the power-law distribution observed in content popularity, which suggests that the most significant insights come from the top rankings. Due to the lack of data for ranks below the 10th, we cannot directly compare the distances calculated using all ranks to those calculated with only the top 10 content. Instead, we assess the robustness of our approach by calculating distances $D_{A,B}$ using only the top 5 ranked contents and comparing this order with the original distances $D_{A,B}$. If these two sets of distances maintain a linear relationship, it confirms that analyzing the top 5 contents is sufficient, thereby supporting the validity of focusing on the top 10 list [^3] For detailed analyses, please refer to the supplementary material.. Another concern arises from the fact that the release dates of content can vary between countries, and in some cases, certain content may be available in one country but not in the other. To verify the validity of our methodology, we recalculated distances using only content that appeared in the top 10 on the same date in both countries. This approach also yielded a strong positive correlation with the distances calculated using our original method, as detailed in the supplementary information. Therefore, we are confident in the validity of our analysis results despite the limitations of approximating countries as points in a hyper-dimensional space and using Euclidean distance to represent differences in VOD consumption across countries.
@@ 2.3 NSC Network and Grouping
Research on analyzing digital consumption data, such as that from Netflix, through network-based approaches has been extensively studied in recent years.[4, 20, 1, 21, 12] In our study, we construct a consumption network where the nodes represent countries and the links between them are weighted by the similarity in content consumption between those countries. Given the distances $D_{A,B}$ between each pair of countries, A and B, we define the weights $W_{A,B}$ as follows:
$W_{A,B}=\frac{D_{\text{min}}-D_{A,B}}{D_{\text{max}}-D_{\text{min}}}$ (5)
where $D_{\text{max}}$ and $D_{\text{min}}$ are the maximum and minimum values of $D_{i,j}$ for all pairs of countries $(i,j)$ with $i\neq j$. Through this formula, the link between the two countries that are furthest apart in the hyperspace is assigned a weight of 0, indicating the lowest similarity, while the link between the closest countries receives a weight of 1, indicating the highest similarity.
With this network, we classify the countries into NSC groups using the Louvain algorithm,[9] a method widely applied across various fields for community detection.[8, 23, 18, 7, 11, 16] The Louvain algorithm detects communities by maximizing the modularity $Q$, which is defined as
$Q=\frac{1}{2m}\sum_{AB}\left[W_{A,B}-\frac{k_{A}k_{B}}{2m}\right]\delta(c_{A},c_{B})$ (6)
@@ 3.1 VOD similar consumption groups
Figure 1 shows the countries grouped by the Louvain algorithm from the data of 1,769 pieces of content, indicated by colors. In all 1,000 simulations, the countries of 71 nations were classified into three groups, and the countries belonging to each group were always consistent except for one country. Looking at the countries in each group, it can be confirmed that they are classified into three VOD consumption groups [^4] North America and Pan-Europe Group: Australia, Austria, Belgium, Bulgaria, Canada, Croatia, Cyprus, Czech Republic, Denmark, Estonia, Finland, France, Germany, Greece, Hungary, Iceland, Ireland, Israel, Italy, Latvia, Lithuania, Netherlands, New Zealand, Norway, Poland, Portugal, Romania, Slovakia, Slovenia, South Africa, Sweden, Switzerland, Ukraine, United Kingdom, United State
@@
The only exception country was Turkiye. Out of 1,000 instances, Turkiye was classified into the Asia and Middle East group 749 times, and into the North America and Pan-Europe group 251 times. It seems to reflect the national character that is geographically on the border of Europe and Asia and has a strong integrative character culturally. In fact, Turkish TV exerts influence in a variety of markets from the Middle East to Europe and Central Asia[17]. It can be inferred that global VOD consumption is strongly influenced by geographical, linguistic, and historical factors.
@@
Figure 2: Top 5 content producing countries and their scores within each group Note. The background color of the graph represents the group being averaged, and the color of the bars represents the group to which the production country belongs. ‘US’ stands for the Unites States, ‘KR’ for South Korea, ‘CO’ for Colombia, ‘UK’ for the United Kingdom, ‘ES’ for Spain, ‘FR’ for France, ‘JP’ for Japan, ‘TR’ for Turkiye, and ‘MX’ for Mexico.
@@
The collected data comprised a total of 1,769 pieces of content, which were produced by 49 countries. To understand how much content produced by individual producing countries was consumed by each country, we considered the sum of scores of the contents produced by a specific producing country as the score for that producing country. In other words, the score for the United States was the sum of scores of 528 pieces of content, and the score for South Korea was the sum of scores of 203 pieces of content. Since each country has contents ranking from 1st to 10th place daily, converting these to scores and summing them up gives 55 points. That is, if all content from 1st to 10th place in a particular country on a particular date was produced by the United States, then the score for the United States in that country on that date would be 55 points. We calculated the average scores for 71 countries over 822 days. The top 5 producing countries with the highest average scores are shown in Fig. 2-(a). Figure 2-(b),(c) and (d) show the results of calculating the average scores for each of three VOD similar consumption groups. Figure 2-(a) represents the average scores over 822 days for all 71 countries, Fig. 2-(b) for the 35 countries in the North America and Pan-Europe group, Fig. 2-(c) for the 18 countries in the Asia and Middle East group, and Fig. 2-(d) for the 18 countries in the Central and South America group. The background color of the graph represents the group averaged, and the color of the bars represents the group the producing country belongs to.
@@
The United States, which ranks first among all countries (see Fig. 2-(a)), occupies high ranks of first, second, and first respectively in the three groups (each in Fig. 2-(b), (c), (d)). South Korea and Spain, which rank second and fifth overall, are also in the Top 5 across all three groups, indicating that content from the United States, South Korea, and Spain consistently receives global approval as a constant variable. In contrast, Colombia and the United Kingdom, which rank third and fourth overall, are localized to the North America and Pan-Europe group and the Central and South America group, respectively, unlike the aforementioned three countries.
@@
Besides the constant variables of the United States, South Korea, and Spain, the countries that rank highly within each group include the United Kingdom and France in the North America and Pan-Europe group, Japan and Turkiye in the Asia and Middle East group, and Colombia and Mexico in the Central and South America group. Interestingly, all of these are countries whose background color and bar color match, that is, they belong to their respective groups. This shows that countries tend to consume a lot of content from countries within the same group. Ultimately, it can be seen that in global VOD consumption, the United States exercises a strong content dominance, South Korea and Spain have mid-range dominance, and a few countries in each consumption group exert a lower level of dominance.

### v3

@@ Abstract
This study explores global video on demand content consumption patterns through a network-based approach. We used Netflix’s ‘TV-shows’ ranking data, spanning 822 days across 71 countries, to construct a network where countries are represented as nodes and consumption similarities are reflected as link weights. By applying the Louvain algorithm, we identified three distinct consumption groups, ‘North America and Pan-Europe’, ‘Asia and Middle East’, and ‘Central and South America group’. These groups align closely with geographic, historical, and linguistic divisions, despite no predefined grouping criteria. Notably, Turkiye, often considered a cultural and regional crossroads, exhibited some classification ambiguity but was ultimately grouped with Asia and Middle East. Our findings also show that the United States accounts for the largest share of content consumption across all groups, while South Korean content, particularly after the success of “Squid Game” in 2021, has gained and maintained popularity in Asia, the Middle East, and Latin America. This study, based on data, demonstrates that deep-seated cultural histories continue to shape global consumption patterns, even amidst rapid changes in media platforms and content production dynamics.
@@ 1 Introduction
Despite ongoing trends of de-globalization in politics and the economy, culture continues to globalize at an unprecedented pace. Video on Demand (VOD) services, such as Netflix, have become major drivers of transnational popular culture. While traditional global media consumption has often followed patterns shaped by cultural proximity[1], viewing behaviors on VOD platforms may be influenced more strongly by geographic proximity. Indeed, Netflix has managed its market by dividing it into four geographic cultural clusters, Europe-Middle East-Africa, Asia-Pacific, The United States-Canada, Latin America.
@@
In spite of interest in VOD acceptance[2], empirical research became possible only after Netflix’s TOP 10 consumption data was made public in 2020. Lotz[3] has analyzed the frequency of overlap in the preferred Netflix Top 10 titles by major countries and argued that global consumption of Netflix consists of large, consistent markets such as English-speaking and Spanish-speaking regions, and small variable markets such as geographically close Korea, Japan, Taiwan, and the unique market of India. Jang et al.[4] have identified eight groups that consume Netflix similarly based on Netflix’s Top 10 consumption data, discovering that they are differentiated by geographical distance, linguistic similarity, and cultural differences.
@@
The concepts of geo-cultural and geo-linguistic markets suggest that global VOD services are consumed differently across regions and languages. Netflix’s 2021 diversity report emphasizes producing content that reflects regional cultures and inclusivity in race and gender[5], a strategy described as a deliberate effort to engage diverse audiences[6, 7]. Yet, important questions remain. These include inquiries into how global VOD consumption is defined and separated, which nations are pivotal in this segmentation, and beyond it. It is also essential to persistently monitor the evolving patterns of consumption among similar viewer groups over time with longitudinal data. What distinct viewer groups emerge from the complex interactions of people connected to VOD platforms worldwide?
@@
This study aims to examine how global VOD is segmented into national and regional markets. These divisions are shaped not only by country borders but also by broader cultural and linguistic areas that often transcend national boundaries. Essentially, we examine how the consumption of VOD services is influenced by and caters to these diverse geo-cultural segments. To give a preliminary report, we found that global VOD content is being consumed in three distinct geo-linguistic consumer groups worldwide. Some countries are located on the boundary of these consumption groups, with Turkiye being a prime example. Due to its unique historical and geographical location straddling the East and West, Turkiye exemplifies a nation that could fall into more than one of these consumption categories.
This may suggest the possibility that global audiences’ choices in VOD consumption are influenced not only by contemporary trends but also by historical and cultural dynamics. The trajectory of VOD media consumption may be seen as reflecting a “long future”, a term used here to highlight the enduring influence of historical developments on present-day cultural practices. Although not formally defined in this context, the phrase echoes Raymond Williams’s idea of a “long revolution”[8] in cultural evolution, which he distinguished from more abrupt transformations like the Industrial or Civil Revolutions. Williams argued that the cultural revolution unfolded gradually over centuries, culminating in the rise of mass media during the 20th century, long after the expansion of literacy and public readership.
In this light, today’s global VOD consumption patterns may represent a continuation of that cultural trajectory, where historical legacies still influence emerging forms of media engagement. This notion of a long cultural momentum underlines the interplay between global media flows and persistent local currents.
@@ 2.1 Data
We analyzed content lists of NS rankings, spanning from the 1st to the 10th place, collected from Flixpatrol over a period of $N_{\text{day}}=822$ days [^1] July 1, 2020 to September 30, 2022 and across $N_{\text{coun}}=71$ countries [^2] The 71 countries mentioned are the sum of the members of the three groups listed in footnote 4.. If each content appeared in the list only once across the period, the total number of content entries would be $822\times 71\times 10=583,620$. However, since most content appears in the list across multiple days and in different countries, the actual number of unique content entries was reduced to $N_{\text{cont}}=1769$.
@@
$G_{A}=\{g_{A}^{1,1},g_{A}^{1,2},\ldots,g_{A}^{2,1},\ldots,g_{A}^{N_{\text{day}},N_{\text{cont}}}\}$ (1)
@@
As equation (2) indicates, the 1st place ranking is converted to 10 points, the 10th place to 1 point, and rankings beyond the 10th place are assigned 0 points. By converting ranking data into score data, we can now represent each country as a point in a high-dimensional space of $822\times 1769=1,454,118$ dimensions. The coordinates of a country in this space are given by its score data. For example, the position of country $A$ in this hyperspace is represented by the set of scores,
$S_{A}$ $=$ $\{s_{A}^{1,1},s_{A}^{1,2},\ldots,s_{A}^{2,1},\ldots,s_{A}^{822,1769}\}.$ (3)
Within this coordinate system, we define the distance between two countries as the Euclidean distance in hyperspace, as detailed below.
@@ 2.2 NSC distance between countries
We aim to define the ‘NSC distance $D_{AB}$’ between two countries as a measure of the difference in their content consumption patterns throughout the entire analysis period. For computational convenience, we first define the daily distance $dd_{AB,d}$ between countries $A$ and $B$ on day $d$, which is based on differences in their ranked contents.
$dd_{AB,d}=\sqrt{\sum_{c=1}^{N_{\text{cont}}}{\left(s_{A}^{d,c}-s_{B}^{d,c}\right)^{2}}}$ (4)
The overall NSC distance $D_{AB}$ is then obtained by summing these daily distances over all days and taking the square root.
$D_{\!AB}=\sqrt{\sum_{d=1}^{N_{\text{day}}}dd_{AB,d}^{2}}=\sqrt{\sum_{d=1}^{N_{\text{day}}}\sum_{c=1}^{N_{\text{cont}}}\left(s_{A}^{d,c}-s_{B}^{d,c}\right)^{2}}$ (5)
@@
In the summation, the terms $\left(s_{A}^{d,c}-s_{B}^{d,c}\right)^{2}$ calculate the squared difference in scores between countries $A$ and $B$ for each piece of content $c$ on day $d$. This approach ensures that all differences are positive and gives more weight to larger discrepancies, regardless of which country had the higher score. Previous research[4] has focused on whether contents were in the top 10 list or not, without considering their specific rankings. We argue that taking ranking scores into account is crucial because the disparity between the 1st and 10th ranks is often more significant than between listed and unlisted content. This is particularly true given the power-law like distribution observed in the popularity of cultural contents[10, 11, 12] which suggests that higher-ranked content has disproportionately greater popularity and influence.
Although we cannot directly analyze the viewer distribution by rank due to the lack of access to actual viewership data per show, we analyzed a related quantity — the total score $s_{{}_{\!T}}^{c}$ for each of the $N_{\text{cont}}$ contents — defined as
$s_{{}_{\!T}}^{c}$ $=$ $\sum_{k=1}^{N_{\text{coun}}}\sum_{d=1}^{N_{\text{day}}}s_{k}^{d,c},$ (6)
which sums the ranking scores of a given content $c$ across all countries and all days in the dataset. We then ranked the contents by their total scores and plotted the distribution, which is shown in Fig. A1 in the Appendix. Although the distribution does not follow power-law in the entire region, but it clearly shows that a few contents received extremely high scores, while most received very low scores. The gap between the most popular and the least popular content is very large, by a factor of about $10^{5}$ (five orders of magnitude). This heavy-tailed pattern supports our assumption that higher-ranked shows tend to attract many more viewers than lower-ranked ones.
@@
In equation (4), the summation index goes from 1 to $N_{\text{cont}}=1769$, but in practice, we only add up at most 20 terms that are not zero. For example, if content $c$ is not in the top 10 for country $A$, then the corresponding term $s_{A}^{d,c}$ is zero, and the same rule applies to country $B$. This implies that the only terms you need to consider are those where the content $c$ is in the top 10 for country $A$ or $B$. All other terms will be zero and can be ignored in the sum.
@@ 2.3 NSC Network and Grouping
Research on analyzing digital consumption data, such as that from Netflix, through network-based approaches has been extensively studied in recent years[13, 14, 15, 16]. In our study, we construct a consumption network where the nodes represent countries and the links between them are weighted by the similarity in content consumption between those countries. Given the distances $D_{\!AB}$ between each pair of countries, $A$ and $B$, we define the weights $W_{\!AB}$ as follows,
$W_{\!AB}=\frac{D_{\text{max}}-D_{\!AB}}{D_{\text{max}}-D_{\text{min}}}$ (7)
where $D_{\text{max}}$ and $D_{\text{min}}$ are the maximum and minimum values of $D_{ij}$ for all pairs of countries $(i,j)$ with $i\neq j$. Through this formula, the link between the two countries that are furthest apart in the hyperspace is assigned a weight of 0, indicating the lowest similarity, while the link between the closest countries receives a weight of 1, indicating the highest similarity.
Some may argue that this weight of the link between two countries may not accurately capture the closeness in actual NSC patterns between two countries due to the limitations set by equation (5). Specifically, one concern arises because the calculation considers only content ranked in the top 10, excluding any data on ranks below the 10th, which could potentially distort the measurement of consumption differences. However, we believe this limitation is negligible due to the power-law distribution observed in content popularity[10, 11, 12], which suggests that the most significant insights come from the top rankings.
Due to the lack of data for ranks below the 10th, we cannot directly compare the weights calculated using all ranks to those calculated with only the top 10 content. Instead, we assess the robustness of our approach by calculating weights $W_{\!AB}^{(5)}$ using only the top 5 ranked contents and comparing this order with the original weights $W_{\!AB}$. If these two sets of weights maintain a linear relationship, it confirms that analyzing the top 5 contents is sufficient, thereby supporting the validity of focusing on the top 10 list.
In Appendix, we show a scatter plot between two variables $W_{\!AB}^{(5)}$ and $W_{\!AB}$ in Fig. A2(a). One can clearly see the linear dependence of these two weights.
Another concern arises from the fact that the release dates of content can vary between countries, and in some cases, certain content may be available in one country but not in the other. To verify the validity of our methodology, we recalculated distances using only content that appeared in the top 10 on the same date in both countries. This approach also yielded a strong positive correlation with the distances calculated using our original method, as detailed in the Appendix. Therefore, we are confident in the validity of our analysis results despite the limitations of approximating countries as points in a hyper-dimensional space and using Euclidean distance to represent differences in VOD consumption across countries.
With this network, we classify the countries into NSC groups using the Louvain algorithm[9], a method widely applied across various fields for community detection[17, 18, 19, 20, 21, 22]. The Louvain algorithm detects communities by maximizing the modularity $Q$, which is defined as
$Q=\frac{1}{2m}\sum_{AB}\left[W_{\!AB}-\frac{k_{A}k_{B}}{2m}\right]\delta(c_{A},c_{B})$ (8)
@@ 3.1 VOD similar consumption groups
Figure 1 shows the countries grouped by the Louvain algorithm from the data of 1,769 pieces of content, indicated by colors. In all 1,000 simulations, the countries of 71 nations were classified into three groups, and the countries belonging to each group were always consistent except for one country. Looking at the countries in each group, it can be confirmed that they are classified into three VOD consumption groups [^3] North America and Pan-Europe Group: Australia, Austria, Belgium, Bulgaria, Canada, Croatia, Cyprus, Czech Republic, Denmark, Estonia, Finland, France, Germany, Greece, Hungary, Iceland, Ireland, Israel, Italy, Latvia, Lithuania, Netherlands, New Zealand, Norway, Poland, Portugal, Romania, Slovakia, Slovenia, South Africa, Sweden, Switzerland, Ukraine, United Kingdom, United States
@@
The only exception was Turkiye. Out of 1,000 instances, Turkiye was classified into the Asia and Middle East group 749 times, and into the North America and Pan-Europe group 251 times. This reflects its geographical position at the border of Europe and Asia and its culturally integrative character. The fluctuation illustrates cultural hybridization, where overlapping ties lead to community inconsistency in Turkiye’s group assignment. Turkish TV also exerts influence in a variety of markets from the Middle East to Europe and Central Asia[23]. Overall, global VOD consumption appears to be strongly associated with geographical, linguistic, and historical factors.
To address the robustness of group detection, we examined the distribution of the modularity score $Q$ across the 1,000 runs. When Turkiye was assigned to the Asia and Middle East group, the modularity was $Q=0.12292$, and when assigned to the North America and Pan-Europe group, it was $Q=0.12299$, showing no significant difference. Although these $Q$ values are below 0.2, a threshold often cited as indicating weak community structure, this outcome arises from the all-to-all nature of our network, which tends to lower absolute $Q$ values. To assess the meaningfulness of the grouping, we generated random groupings with the same group sizes (35, 18, and 18 countries) and calculated $Q$ 1,000 times. The random groupings yielded an average Q=0.009193Q=-0.009193 with a standard deviation of $0.003706$, giving a $z$-score of $35.64$ for our observed result. This large positive $z$ demonstrates that, despite the low absolute $Q$ values, the detected grouping is highly significant compared with random partitions.
We also note that our results are not sensitive to the specific definition of similarity weights between countries. As shown in Appendix A2, different definitions of similarity weights exhibit a strong linear relationship with each other, and the resulting group classifications remain largely consistent across these alternatives. In particular, when cosine similarity is used, which depends only on the angle between country vectors in the hyperspace rather than their Euclidean distance, the analysis produces the same three VOD similarity groups with identical group memberships as those obtained by our original method. This indicates that the observed group structure reflects stable patterns of VOD consumption and is not strongly affected by the choice of similarity metric.
@@
Figure 2: Top 5 content producing countries and their scores within each group. The background color of the graph represents the group being averaged, and the color of the hatched bars represents the group that the producing country belongs to. ‘US’ stands for the United States, ‘KR’ for South Korea, ‘CO’ for Colombia, ‘UK’ for the United Kingdom, ‘ES’ for Spain, ‘FR’ for France, ‘JP’ for Japan, ‘TR’ for Turkiye, and ‘MX’ for Mexico.
The dataset contains a total of 1,769 titles produced across 49 countries. The production country of each title was identified using metadata provided by FlixPatrol, which classifies origin based on publicly available information such as the country of the main production company or the official release country. To evaluate how much content from each producing country was consumed globally, we define a production score for each country. The production score of country $p$ is defined as the sum of the total scores of all contents produced in that country,
$S_{p}$ $=$ $\sum_{c\in C_{p}}s_{T}^{c},$ (9)
where $C_{p}$ is the set of contents produced in country $p$, and $s_{T}^{c}$ is the total score of content $c$ as defined in Eq. (6). For instance, the production score for the United States is obtained by summing the total scores of 528 US-produced titles, while that for South Korea is based on 203 South Korean titles. This measure provides a way to compare the relative global presence of producing countries in terms of VOD consumption.
@@
Since each country has contents ranking from 1st to 10th place daily, converting these to scores and summing them up gives 55 points. That is, if all content from 1st to 10th place in a particular country on a particular date was produced by the United States, then the score for the United States in that country on that date would be 55 points. We calculated the average scores for 71 countries over 822 days. The top 5 producing countries with the highest average scores are shown in Fig. 2-(a). Figure 2-(b),(c) and (d) show the results of calculating the average scores for each of three VOD similar consumption groups. Figure 2-(a) represents the average scores over 822 days for all 71 countries, Fig. 2-(b) for the 35 countries in the North America and Pan-Europe group, Fig. 2-(c) for the 18 countries in the Asia and Middle East group, and Fig. 2-(d) for the 18 countries in the Central and South America group. The background color of the graph shows the group over which the average is taken, while the color of the hatched bars shows the group of the producing country.
@@
The United States, which ranks first among all countries (see Fig. 2-(a)), occupies high ranks of first, second, and first respectively in the three groups (each in Fig. 2-(b), (c), (d)). South Korea and Spain, which rank second and fifth overall, are also in the Top 5 across all three groups, indicating that content from the United States, South Korea, and Spain consistently receives global approval as a constant variable. In contrast, Colombia and the United Kingdom, which rank third and fourth overall, are localized to the Central and South America group and the North America and Pan-Europe group respectively, unlike the aforementioned three countries.
@@
Besides the constant variables of the United States, South Korea, and Spain, the countries that rank highly within each group include the United Kingdom and France in the North America and Pan-Europe group, Japan and Turkiye in the Asia and Middle East group, and Colombia and Mexico in the Central and South America group. Interestingly, all of these are countries whose background color and bar color match, that is, they belong to their respective groups. This shows that countries tend to consume a lot of content from countries within the same group. Ultimately, the United States, South Korea, and Spain are characterized by content that attracts audiences across multiple groups, whereas the content from other countries is primarily consumed within their own group.